Suppose an investment will pay $11,000 in 29 years from now. If you can earn 15.35% interest compounded monthly by depositing your money in a bank, how much should you pay for the investment today?

Round your answer to two decimal places. For example, if your answer is $345.667 round as 345.67 and if your answer is .05718 or 5.718% round as 5.72.
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

11000=P(1+0.1535/12)^(12*29)

P=11000/(1+0.1535/12)^(12*29)

=$11000*0.011994951

which is equal to

=$131.94(Approx).